### v2

# On the Global Optimality of Linear Policies for Sinkhorn Distributionally Robust Linear Quadratic Control

Abstract
The Linear Quadratic Gaussian (LQG) regulator is a cornerstone of optimal control theory, yet its performance can degrade significantly when the noise distributions deviate from the assumed Gaussian model. To address this limitation, this work proposes a distributionally robust generalization of the finite-horizon LQG control problem. Specifically, we assume that the noise distributions are unknown and belong to ambiguity sets defined in terms of a Sinkhorn discrepancy centered at a nominal Gaussian distribution. By deriving novel bounds on this entropy-regularized Wasserstein distance and proving structural and topological properties of the resulting ambiguity sets, we establish global optimality of linear policies for Sinkhorn distributionally robust LQG. Numerical experiments showcase improved distributional robustness of our control policy.
I Introduction
The theory of Linear Quadratic Gaussian (LQG) regulators addresses the fundamental problem of controlling partially-observed linear systems driven by additive Gaussian noise with the objective of minimizing an expected quadratic cost [4]. This problem admits an elegant closed-form solution, combining a Kalman filter with a linear state-feedback controller, and has found application in a variety of domains ranging from engineering to economics and computer science.
In the presence of model misspecifications, however, the LQG solution can be extremely fragile [9]. Classical $\mathcal{H}_{\infty}$ control [32] addresses this concern by shifting from a stochastic to an adversarial uncertainty model and minimizing the worst-case cost across bounded-energy disturbances. While provably robust, $\mathcal{H}_{\infty}$ methods tend to be overly conservative in practice as they optimize for the least favorable uncertainty realization. Motivated by this observation, several approaches have been proposed to balance nominal performance and robustness, including mixed $\mathcal{H}_{2}/\mathcal{H}_{\infty}$ formulations [3, 8], risk-sensitive control [17], and regret minimization methods [24, 15, 23].
Among these approaches, recent work on distributionally robust (DR) control promises to combine the advantages of stochastic and adversarial uncertainty models by robustifying in the probability space. To achieve this, the DR control paradigm considers the problem of minimizing the expected cost under the most averse distribution within a given ambiguity set—a set of distributions that are sufficiently close, in an appropriate sense, to a nominal one. For instance, [29] studied DR control of constrained stochastic systems with ambiguity sets comprising all distributions sharing the same first two moments. To account for full distributional information, the works [25, 10, 11] instead employ ambiguity sets defined in terms of $f$-divergence, whereas [28, 21, 31, 5, 1, 19] rely on optimal transport metrics in light of their proven expressiveness and out-of-distribution guarantees.
Inspired by these results and motivated by the recent application of the Sinkhorn discrepancy for DR learning [30, 14] and control [6], we study a generalization of the finite-horizon LQG control problem, where the noise distributions are unknown and belong to Sinkhorn ambiguity sets centered at nominal Gaussian distributions. Our main contribution is to establish the global optimality of linear policies for the Sinkhorn DR LQG problem. While our proof follows a “sandwich” argument akin to [28], the extension from Wasserstein DR LQG to Sinkhorn DR LQG is far from direct: the introduction of the KL regularization term fundamentally alters the geometry of the ambiguity set [^1] For instance, it is well-known that the worst-case distribution in Wasserstein DR optimization is finitely supported when so is the center distribution [12], while the Sinkhorn ambiguity set with a Gaussian reference measure only contains continuous distributions. and invalidates several key inequalities used in [28]. To recover a comparable optimality result, we develop new tools specific to the Sinkhorn framework. In particular, we first construct a novel lower bound to our DR LQG problem leveraging results from regularized optimal transport [7]; second, we establish a Gelbrich-type inequality for the Sinkhorn discrepancy, which is not available in the literature and is essential for deriving an upper bound to our DR LQG problem; last, we prove convexity and compactness of the resulting entropy-regularized Gelbrich ambiguity set. Taken together, these results allow us to conclude that the Sinkhorn DR LQG admits a Nash equilibrium in the class of linear policies.
Alongside [28, 21, 11, 10], our work contributes to delineating scenarios where, despite the additional complexity introduced by DR formulations, linear feedback policies remain globally optimal for LQG control problems.
Notation: Throughout the paper, we denote the set of (zero-mean) probability distributions supported on a measurable set $\mathcal{Z}$ by $\mathcal{P}(\mathcal{Z})$ (resp. $\mathcal{P}_{0}(\mathcal{Z})$). We write $\mu\ll\nu$ to denote that a measure $\mu$ is absolutely continuous with respect to $\nu$. The convolution product between two probability measures is represented by $\mu*\nu$. For $n\in\mathbb{N}$, we write $[n]$ to denote the set of indices $\{0,\dots,n-1\}$. The space of all $d\times d$ positive (semi)definite matrices is denoted by $\mathbb{S}_{++}^{d}$ (resp. $\mathbb{S}^{d}_{+}$). We denote by $\|\cdot\|$ the Euclidean norm. The determinant of a square matrix $A$ is denoted by $|A|$. Given $A\in\mathbb{S}^{d}$, we denote by $\{\lambda_{i}(A)\}_{i=1}^{d}$ its eigenvalues and let $\lambda_{\text{max}}(A)=\max_{i}\lambda_{i}(A)$.
II Preliminaries
We begin by recalling definitions of discrepancies between probability distributions that will be used throughout the paper.
Definition 1 (KL divergence, [20, Definition 2.8]):
Given $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{d})$ with $\mathbb{P}\ll\mathbb{Q}$, the Kullback-Leibler (KL) divergence between $\mathbb{P}$ and $\mathbb{Q}$ is
$\mathrm{KL}(\mathbb{P}\|\mathbb{Q})=\mathbb{E}_{\mathbb{P}}\left[\log\left(\frac{\mathop{}\!\mathrm{d}\mathbb{P}(x)}{\mathop{}\!\mathrm{d}\mathbb{Q}(x)}\right)\right].$
Definition 2 (Sinkhorn discrepancy, [30, Definition 1]):
Let $\mathbb{P},\mathbb{Q}\in\mathcal{P}(\mathbb{R}^{d})$ and $\mu,\nu$ be reference probability measures over $\mathbb{R}^{d}$ such that $\mathbb{P}\ll\mu$ and $\mathbb{Q}\ll\nu$. For any $\epsilon\geq 0$, the Sinkhorn discrepancy between $\mathbb{P}$ and $\mathbb{Q}$ is defined as
$W_{\epsilon}(\mathbb{P},\mathbb{Q})=\!\inf_{\gamma\in\Gamma(\mathbb{P},\mathbb{Q})}\left\{\mathbb{E}_{\gamma}[\|x-y\|^{2}]+\\ \epsilon\mathrm{KL}(\gamma\|\mu\times\nu)\right\},$ (1)
where $\Gamma(\mathbb{P},\mathbb{Q})$ denotes the set of all couplings $\gamma$ between $\mathbb{P}$ and $\mathbb{Q}$, that is, the set of all joint distributions with marginals $\mathbb{P}$ and $\mathbb{Q}$.
As noted in [30, Remark 2], any choice of $\mathbb{P}\ll\mu$ in (1) is equivalent up to a constant. Hence, as in our DR LQG problem the center distribution is assumed to be known, we let $\mu=\mathbb{P}$ without loss of generality. Aligned with [30, 6], in the following we further assume $\nu\sim\mathcal{N}(0,\Sigma)$; this choice ensures that $W_{\epsilon}(\mathbb{P},\mathbb{Q})$ is finite for any $\mathbb{Q}$ in the ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$, independently of its support.
Remark 1:
Other definitions of Sinkhorn discrepancy have also been considered in the literature [7, 18, 22]. For instance, [7] regularizes the transport cost with the negative differential entropy of the transport plan $\gamma$, while [18, 22] use (1) with $\mu=\mathbb{P}$ and $\nu=\mathbb{Q}$. Crucially, all these definitions lead to the same optimal transport plan $\gamma^{\star}$. In fact, one can observe that
$\mathrm{KL}(\gamma\|\mu\times\nu)=\mathrm{KL}(\gamma\|\mathbb{P}\times\mathbb{Q})+\mathrm{KL}(\mathbb{P}\times\mathbb{Q}\|\mu\times\nu)\,;$
hence, $\mathrm{KL}(\gamma\|\mu\times\nu)$ and $\mathrm{KL}(\gamma\|\mathbb{P}\times\mathbb{Q})$ are equivalent up to a term that is independent of $\gamma$. A similar reasoning applies when the negative differential entropy is used as regularization term [7].
Definition 3:
Let $\nu\sim\mathcal{N}(0,\Sigma)$ where $\Sigma\in\mathbb{S}^{d}_{++}$. The entropy-regularized Gelbrich divergence between two probabilities $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ with covariance matrices $\Sigma_{1},\Sigma_{2}\in\mathbb{S}_{++}^{d}$ is
$G_{\epsilon}(\Sigma_{1},\Sigma_{2})\!$ $=\!\operatorname{Tr}(\Sigma_{1})\!+\!\operatorname{Tr}(\Sigma_{2})\!-\!2\operatorname{Tr}(D_{\epsilon})\!+\!\frac{\epsilon}{2}\operatorname{Tr}\left(\Sigma^{-1}\Sigma_{2}\right)$
$+\frac{\epsilon}{2}\log\frac{|\Sigma|}{|\Sigma_{2}|}\!+\!\frac{\epsilon}{2}\log\left(\left(\frac{2}{\epsilon}\right)^{d}\left|D_{\epsilon}+\frac{\epsilon}{4}I\right|\right),$ (2)
where $D_{\epsilon}=\left(\Sigma_{1}^{\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{\frac{1}{2}}+\frac{\epsilon^{2}}{16}I\right)^{\frac{1}{2}}$.
Unlike (1), Definition 3 only accounts for the covariances of $\mathbb{P}_{1},\mathbb{P}_{2}\in\mathcal{P}_{0}(\mathbb{R}^{d})$. In Section IV, we will exploit connections between (1) and (2) to construct finite-dimensional upper and lower bounds to our Sinkhorn DR LQG problem.
[…]
$\underline{\rho}=\tfrac{\epsilon}{2}\big(\operatorname{Tr}\!\big(\Sigma^{-1}(\hat{\Sigma}+\tfrac{\epsilon}{2}I)\big)-d+\log|\Sigma|-d\log(\tfrac{\epsilon}{2})\big)\,.$ (3)
This value represents the smallest radius $\rho$ such that the ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$ is non-empty. Last, we remark that, when $\epsilon\rightarrow 0$, Definition 3 is equivalent to the squared Gelbrich distance, see [13, Theorem 2.1].
III Problem Formulation
We consider discrete-time linear dynamical systems described by the following state-space equations
$x_{t+1}=A_{t}x_{t}+B_{t}u_{t}+w_{t}\,,\quad y_{t}=C_{t}x_{t}+v_{t}\,,~\forall t\in[T]\,$, (4)
where $x_{t}\in\mathbb{R}^{d}$ denotes the state vector, $u_{t}\in\mathbb{R}^{m}$ the control input, $y_{t}\in\mathbb{R}^{p}$ the output, $w_{t}\in\mathbb{R}^{d}$ the process noise, $v_{t}\in\mathbb{R}^{p}$ the measurement noise, and $T\in\mathbb{N}$ the control horizon.
For ease of presentation, we collect all exogenous random vectors in the variable $\delta=(x_{0},w_{0},\dots,w_{T-1},v_{0},\dots,v_{T-1})$. We assume that all entries of $\delta$ are mutually independent. Differently from classical LQG theory, however, we assume that their true distributions are unknown and belong to a Sinkhorn ambiguity set $\mathcal{S}$ centered at a nominal distribution $\hat{\mathbb{P}}=\hat{\mathbb{P}}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\hat{\mathbb{P}}_{v_{t}}\right)$, with $\hat{\mathbb{P}}_{x_{0}}=\mathcal{N}(0,\hat{X}_{0})$, $\hat{\mathbb{P}}_{w_{t}}=\mathcal{N}(0,\hat{W}_{t})$, and $\hat{\mathbb{P}}_{v_{t}}=\mathcal{N}(0,\hat{V}_{t})$, for all $t\in[T]$. [^2] Following [28, 21, 11, 10], we consider only distributions with zero mean; our results extend to the non-zero mean case with minor modifications. Specifically, for any regularization parameter $\epsilon\geq 0$ and user-defined radii $\rho_{x_{0}}\geq\underline{\rho}_{x_{0}},\rho_{w_{t}}\geq\underline{\rho}_{w_{t}},\rho_{v_{t}}\geq\underline{\rho}_{v_{t}}$, where $\underline{\rho}_{x_{0}},\underline{\rho}_{w_{t}}$, and $\underline{\rho}_{v_{t}}$ are defined according to (3) with $\hat{X}_{0}$, $\hat{W}_{t}$, and $\hat{V}_{t}$ in place of $\hat{\Sigma}$, respectively, we define the ambiguity set $\mathcal{S}$ as $\mathcal{S}_{x_{0}}\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{w_{t}}\right)\otimes\left(\otimes_{t=0}^{T-1}\mathcal{S}_{v_{t}}\right)$ where
$\mathcal{S}_{x_{0}}$ $=\left\{\mathbb{P}_{x_{0}}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\hat{\mathbb{P}}_{x_{0}},\mathbb{P}_{x_{0}})\leq\rho_{x_{0}}\right\}\,$,
$\mathcal{S}_{w_{t}}$ $=\left\{\mathbb{P}_{w_{t}}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\hat{\mathbb{P}}_{w_{t}},\mathbb{P}_{w_{t}})\leq\rho_{w_{t}}\right\}\,$,
$\mathcal{S}_{v_{t}}$ $=\left\{\mathbb{P}_{v_{t}}\in\mathcal{P}_{0}(\mathbb{R}^{p}):W_{\epsilon}(\hat{\mathbb{P}}_{v_{t}},\mathbb{P}_{v_{t}})\leq\rho_{v_{t}}\right\}\,.$
Given a realization $\delta$ of the exogenous vectors and a collection of causal measurable [^3] Throughout the paper, we tacitly assume that the probability space of the exogenous signals is equipped with the standard Borel $\sigma$-algebra. functions $\pi_{t}:\mathbb{R}^{p(t+1)}\rightarrow\mathbb{R}^{m}$ mapping output observations to control inputs as per $u_{t}=\pi_{t}(y_{0:t})$, we define the cost incurred by the policy $\pi=(\pi_{0},\dots,\pi_{T-1})$ as
$J(\pi,\delta)=\sum_{t=0}^{T-1}(x_{t}^{\top}Q_{t}x_{t}+u_{t}^{\top}R_{t}u_{t})+x_{T}^{\top}Q_{T}x_{T}\,$,
where $Q_{t},Q_{T}\in\mathbb{S}^{d}_{+}$ and $R_{t}\in\mathbb{S}^{m}_{++}$ represent state and input weight matrices, respectively. With these definitions in place, we formulate the Sinkhorn DR LQG control problem as
$\inf_{\pi\in\mathcal{U}_{y}}\ \max_{\mathbb{P}\in\mathcal{S}}\ \mathbb{E}_{\mathbb{P}}[J(\pi,\delta)]\,$, (5)
where $\mathcal{U}_{y}$ denotes the set of all feasible control inputs $\bm{u}=(u_{0},\dots,u_{T-1})$. In particular, (5) can be interpreted as a zero-sum game between the control designer, who selects a causal policy to minimize the expected cost, and an adversary, who chooses the noise distributions in $\mathcal{S}$ that maximize such cost.
IV Analysis of the Sinkhorn DR LQG problem
We now present our main results. We first re-parametrize (5) in terms of the purified observations [2, Section 14.4.2]. Then, we construct two auxiliary problems over the class of linear policies that provide lower and upper bounds to (5). Last, using a “sandwich” argument, we show that the optimal value of these two auxiliary problems coincide, implying that (5) admits a Nash equilibrium and that linear policies are globally optimal.
IV-A Purified output re-parametrization
To ease analysis of (5), we rewrite $u_{t}$ in terms of the purified observations instead of the actual observations $y_{0:t}$. To define these new variables, we introduce a noise-free copy of (4) as
$\hat{x}_{t+1}=A_{t}\hat{x}_{t}+B_{t}u_{t},\quad\hat{y}_{t}=C_{t}\hat{x}_{t},~\forall t\in[T],$ (6)
with state $\hat{x}_{t}\in\mathbb{R}^{d}$ and output $\hat{y}_{t}\in\mathbb{R}^{p}$, initialized at $\hat{x}_{0}=0$ and with the same input $u_{t}$ as the original system. We then define the purified output $\eta_{t}$ at time $t$ as $\eta_{t}=y_{t}-\hat{y}_{t}$, and let $\bm{\eta}=(\eta_{0},\dots,\eta_{T-1})$. This representation proves useful because, as shown in [16, Proposition II.1], every measurable function of $y_{0},\dots,y_{t}$ can be equivalently expressed as a measurable function of $\eta_{0},\dots,\eta_{t}$, and vice versa—yet differently from $\mathbf{y}=(y_{0},\dots,y_{T-1})$, $\bm{\eta}$ is independent of the inputs. In fact, using (4) and (6) recursively, one can show that $\bm{\eta}$ only depends on the exogenous vectors $\mathbf{w}=(x_{0},w_{0},\dots,w_{T-1})$ and $\mathbf{v}=(v_{0},\dots,v_{T-1})$. In particular, it holds that $\bf\bm{\eta}=Dw+v$, where $\bf D=CG$ and $\bf C$ and $\bf G$ are matrices defined in the Appendix.
In light of this, we have that $\mathcal{U}_{y}=\mathcal{U}_{\eta}$, where $\mathcal{U}_{\eta}$ denotes the set of input sequences $\mathbf{u}=(u_{0},\dots,u_{T-1})$ for which there exist measurable functions $\tilde{\pi}_{t}:\mathbb{R}^{p(t+1)}\to\mathbb{R}^{m}$ satisfying $u_{t}=\tilde{\pi}_{t}(\eta_{0:t})$. Hence, we equivalently rewrite problem (5) as
$p^{\star}=\left\{\begin{aligned} &\min_{\bf u}\ \max_{\mathbb{P}\in\mathcal{S}}\ \mathbb{E}_{\mathbb{P}}\left[\bf u^{\top}Ru+x^{\top}Qx\right]\\ &\text{ s.t. }\ \mathbf{u}\in\mathcal{U}_{\eta},\quad\bf x=Hu+Gw,\end{aligned}\right$. (7)
where $\mathbf{x}=(x_{0},\dots,x_{T})$ and $\bf R,Q,H$ are suitable matrices defined in the Appendix.
With the objective of establishing the existence of a Nash equilibrium, we also define the dual problem of (7) as
$d^{\star}=\left\{\begin{aligned} &\max_{\mathbb{P}\in\mathcal{S}}\ \min_{\bf u}\ \mathbb{E}_{\mathbb{P}}\left[\bf u^{\top}Ru+x^{\top}Qx\right]\\ &\text{ s.t. }\ \mathbf{u}\in\mathcal{U}_{\eta},\quad\bf x=Hu+Gw.\end{aligned}\right$. (8)
Classical min-max inequality states that $d^{\star}\leq p^{\star}$. In the next sections, we prove that such relation actually holds with equality—despite the fact that $\mathcal{U}_{\eta}$ is an infinite-dimensional function space and $\mathcal{S}$ is an infinite-dimensional set of non-parametric probability distributions. In particular, our analysis reveals that there exists a Nash equilibrium of the zero-sum game (5) in the form $(\mathbf{u}^{\star},\mathbb{P}^{\star})$, where $\bf u^{\star}=U^{\star}\bm{\eta}+q^{\star}$ is an affine policy of $\bm{\eta}$ and $\mathbb{P}^{\star}$ is Gaussian. To do so, we first construct a lower-bound to (8) and an upper-bound to (7), and then show that their optimal values coincide.
IV-B Construction of a lower bound for $d^{\star}$
To derive a lower bound for the dual problem (8), we restrict our attention to the set $\mathcal{S}_{\mathcal{N}}\subseteq\mathcal{S}$ of Gaussian distributions contained in $\mathcal{S}$. This leads us to the optimization problem
$\underline{d}^{\star}=\left\{\begin{aligned} &\max_{\mathbb{P}\in\mathcal{S_{N}}}\ \min_{\bf u}\ \mathbb{E}_{\mathbb{P}}\left[\bf u^{\top}Ru+x^{\top}Qx\right]\\ &\text{ s.t. }\ \mathbf{u}\in\mathcal{U}_{\eta},\quad\bf x=Hu+Gw.\end{aligned}\right$. (9)
Compared to (8), in (9) we restricted the feasible set in the outer maximization. Hence, we have $\underline{d}^{\star}\leq d^{\star}$. On the other hand, (9) is still an infinite-dimensional optimization problem. In the remainder of this section, we show that this problem can be reformulated as a finite-dimensional one by exploiting known closed-form expressions for the Sinkhorn discrepancy between two normal distributions.
Proposition 1:
[…]
$X_{\epsilon}=\Sigma_{1}^{-\frac{1}{2}}\left(\Sigma_{1}^{\frac{1}{2}}\Sigma_{2}\Sigma_{1}^{\frac{1}{2}}+\frac{\epsilon^{2}}{16}I\right)^{\frac{1}{2}}\Sigma_{1}^{-\frac{1}{2}}-\frac{\epsilon}{4}\Sigma_{1}^{-1}\,.$
Moreover, it holds that $W_{\epsilon}(\mathbb{P}_{1},\mathbb{P}_{2})=G_{\epsilon}(\Sigma_{1},\Sigma_{2})$, that is, the Sinkhorn divergence coincides with the entropy-regularized Gelbrich divergence.
As observed in Remark 1, regularizing (1) with the negative differential entropy or a KL term leads to the same optimal transport plan. Hence, the expression for $\gamma_{0}$ follows from [7, Theorem 2.2]. Substituting this optimal coupling in the definition of Sinkhorn discrepancy in (1), we obtain
$\mathbb{E}_{\gamma_{0}}[\|x-y\|^{2}]$ $=\operatorname{Tr}(\Sigma_{1})+\operatorname{Tr}(\Sigma_{2})-2\operatorname{Tr}(\Sigma_{1}X_{\epsilon})$
$\mathrm{KL}(\gamma_{0}\|\mathbb{P}_{1}\times\nu)$ $=\frac{1}{2}\left[\operatorname{Tr}(\Sigma^{-1}\Sigma_{2})-d+\log\frac{|\Sigma|}{|\Sigma_{2}|}+\right$.
$\left.+\log\left(\left(\frac{2}{\epsilon}\right)^{d}\left|D_{\epsilon}+\frac{\epsilon}{4}I\right|\right)\right].$
By inspection, combining the expressions above as per (1) yields (2), which concludes the proof. ∎ Let us define the matrices $\mathbf{M}=\mathbf{R}+\mathbf{H}^{\top}\mathbf{QH}\in\mathbb{R}^{mT}$, $\mathbf{F}_{1}=\mathbf{D}^{\top}\mathbf{U}^{\top}\mathbf{R}\mathbf{U}\mathbf{D}+(\mathbf{HUD}+\mathbf{G})^{\top}\!\mathbf{Q}(\mathbf{HUD}+\mathbf{G})\in\mathbb{S}_{+}^{d(T+1)}$, and $\mathbf{F}_{2}=\mathbf{U}^{\top}\mathbf{R}\mathbf{U}+\mathbf{U}^{\top}\mathbf{H}^{\top}\mathbf{QHU}\in\mathbb{S}_{+}^{pT}$ for brevity. Moreover, we denote the set of causal feedback matrices by $\mathcal{U}^{\text{lin}}$. With this notation in place, we are now ready to reformulate (9) as a finite-dimensional optimization problem.
Proposition 2:
The lower bound (9) to the dual problem (8) is equivalent to the finite-dimensional optimization problem
$\underline{d}^{\star}=\max_{\begin{subarray}{c}\mathbf{W}\in\mathcal{G}_{W}\\ \mathbf{V}\in\mathcal{G}_{V}\end{subarray}}\ \min_{\begin{subarray}{c}\mathbf{U}\in\mathcal{U}^{\text{lin}}\\ \mathbf{q}\in\mathbb{R}^{mT}\end{subarray}}\ \operatorname{Tr}\left(\mathbf{F}_{1}\mathbf{W}+\mathbf{F}_{2}\mathbf{V}\right)+\mathbf{q}^{\top}\mathbf{M}\mathbf{q},$ (10)
where $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ depend on $\bf U$, and the finite-dimensional sets $\mathcal{G}_{W}$ and $\mathcal{G}_{V}$ are defined as
$\mathcal{G}_{W}$ $=\big\{\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)}:\ \mathbf{W}=\operatorname{diag}(X_{0},W_{0},\ldots,W_{T{-}1}),$
$\quad\quad X_{0}\in\mathbb{S}_{++}^{d},W_{t}\in\mathbb{S}_{++}^{d},\ G_{\epsilon}(\hat{X}_{0},X_{0})\leq\rho_{x_{0}},$
$\quad\quad G_{\epsilon}(\hat{W}_{t},W_{t})\leq\rho_{w_{t}}\ \forall t\in[T]\big\},$ (11)
$\mathcal{G}_{V}$ $=\big\{\mathbf{V}\in\mathbb{S}_{++}^{pT}:\ \mathbf{V}=\operatorname{diag}(V_{0},\ldots,V_{T{-}1}),\ V_{t}\in\mathbb{S}_{++}^{p},$
$\quad\quad G_{\epsilon}(\hat{V}_{t},V_{t})\leq\rho_{v_{t}}\ \forall t\in[T]\big\}.$ (12)
We first observe that, for any fixed $\mathbb{P}\in\mathcal{S_{N}}$, the inner minimization in (9) constitutes a standard LQG problem, for which linear policies are globally optimal [4]. Hence, as discussed in Section IV-A, we restrict the inner minimization in (9) to policies of the form $\bf u=U\bm{\eta}+q$, where $\mathbf{q}\in\mathbb{R}^{mT}$ and $\bf U\in\mathcal{U}^{\text{lin}}$, without loss of generality.
Then, we note that, by Proposition 1, the set $\mathcal{S_{N}}$ is equivalent to the entropy-regularized Gelbrich set
$\mathcal{G}=\mathcal{G}_{x_{0}}\otimes\Big(\otimes_{t=0}^{T-1}\mathcal{G}_{w_{t}}\Big)\otimes\Big(\otimes_{t=0}^{T-1}\mathcal{G}_{v_{t}}\Big)\,$, (13)
where each component $\mathcal{G}_{x_{0}}$, $\mathcal{G}_{w_{t}}$, and $\mathcal{G}_{v_{t}}$ is defined by
$\mathcal{G}_{x_{0}}\!$ $=\!\big\{\mathbb{P}_{x_{0}}\!\in\!\mathcal{P}_{0}(\mathbb{R}^{d})\!:\!\mathbb{E}_{\mathbb{P}}[x_{0}x_{0}^{\top}]=X_{0},G_{\epsilon}(\hat{X}_{0},X_{0})\!\leq\!\rho_{x_{0}}\big\},$
$\mathcal{G}_{w_{t}}\!$ $=\!\big\{\mathbb{P}_{w_{t}}\!\in\!\mathcal{P}_{0}(\mathbb{R}^{d})\!:\!\mathbb{E}_{\mathbb{P}}[w_{t}w_{t}^{\top}]=W_{t},G_{\epsilon}(\hat{W}_{t},W_{t})\!\leq\!\rho_{w_{t}}\big\},$
$\mathcal{G}_{v_{t}}\!$ $=\!\big\{\mathbb{P}_{v_{t}}\!\in\!\mathcal{P}_{0}(\mathbb{R}^{p})\!:\!\mathbb{E}_{\mathbb{P}}[v_{t}v_{t}^{\top}]=V_{t},G_{\epsilon}(\hat{V}_{t},V_{t})\!\leq\!\rho_{v_{t}}\big\}.$
Combining these observations, we equivalently rewrite (9) as
$\max_{\mathbb{P}\in\mathcal{G}}\ \min_{\bf U,q}\ \mathbb{E}_{\mathbb{P}}\left[\bf u^{\top}Ru+x^{\top}Qx\right]$
$\text{ s.t. }\,\,\mathbf{U}\in\mathcal{U}^{\text{lin}},\ \bf{u=U(Dw+v)+q},~~\bf{x=Hu+Gw}.$
Following the same argument as [28, Proposition 3.2], the proof is concluded by rewriting the expectation of a quadratic form as a trace and by replacing the ambiguity set $\mathcal{G}$ by (11) and (12). ∎
IV-C Construction of an upper bound for $p^{\star}$
To derive an upper bound for $p^{\star}$, we restrict our attention to linear policies in $\mathcal{U}_{\eta}$ and appropriately enlarge the ambiguity set $\mathcal{S}$. This construction relies on the following result, which provides a lower bound to the Sinkhorn discrepancy between two distributions when one of them is Gaussian.
Proposition 3:
Let $\mathbb{P}\sim\mathcal{N}(0,\Sigma_{1})$, $\Sigma_{1}\in\mathbb{S}_{++}^{d}$, and $\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d})$ be a distribution with covariance $\Sigma_{2}\in\mathbb{S}_{++}^{d}$. Then, it holds that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$.
Recall the definition of the joint probability distribution $\gamma_{0}$ given in Proposition 1. For any $\gamma\in\Gamma(\mathbb{P},\mathbb{Q})$, the objective in (1) can be rewritten as
$\epsilon\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\left(\frac{\mathop{}\!\mathrm{d}\gamma(x,y)e^{\frac{\|x-y\|^{2}}{\epsilon}}}{\mathop{}\!\mathrm{d}\mathbb{P}(x)\mathop{}\!\mathrm{d}\nu(y)}\right)\mathop{}\!\mathrm{d}\gamma(x,y)$
$=\epsilon\mathrm{KL}(\gamma\|\gamma_{0})+\epsilon\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\left(\frac{\mathop{}\!\mathrm{d}\gamma_{0}(x,y)e^{\frac{\|x-y\|^{2}}{\epsilon}}}{\mathop{}\!\mathrm{d}\mathbb{P}(x)\mathop{}\!\mathrm{d}\nu(y)}\right)\mathop{}\!\mathrm{d}\gamma(x,y)$
$=\epsilon\mathrm{KL}(\gamma\|\gamma_{0})+\epsilon\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\log\left(\mathop{}\!\mathrm{d}\gamma_{0}(x,y)e^{\frac{\|x-y\|^{2}}{\epsilon}}\right)\mathop{}\!\mathrm{d}\gamma(x,y)$
$\quad-\epsilon\int_{\mathbb{R}^{d}}\log\left(\mathop{}\!\mathrm{d}\mathbb{P}(x)\right)\mathop{}\!\mathrm{d}\mathbb{P}(x)-\epsilon\int_{\mathbb{R}^{d}}\log\left(\mathop{}\!\mathrm{d}\nu(y)\right)\mathop{}\!\mathrm{d}\mathbb{Q}(y)$
$=\underbrace{\operatorname{Tr}(\Sigma_{1}+\Sigma_{2})\!-\!2\operatorname{Tr}(\Sigma_{1}X_{\epsilon})\!-\!\tfrac{\epsilon}{2}\log\!\left((2\pi e)^{2d}\!\left(\tfrac{\epsilon}{2}\right)^{d}\!|\Sigma_{1}X_{\epsilon}|\right)}_{(\spadesuit)}$
$+\!\underbrace{\tfrac{\epsilon}{2}\log\!\left((2\pi e)^{d}|\Sigma_{1}|\right)}_{(\clubsuit)}\!+\!\underbrace{\tfrac{\epsilon}{2}\left(\operatorname{Tr}(\Sigma^{-1}\Sigma_{2})\!-\!d\!+\!\log\!\left((2\pi e)^{d}|\Sigma|\right)\right)}_{(\diamondsuit)}$
$+\epsilon\mathrm{KL}(\gamma\|\gamma_{0}),$
where $(\spadesuit)$ is obtained computing the integral with the explicit expression of the density of $\gamma_{0}$ as per Proposition 1, $(\clubsuit)$ is the differential entropy of the Gaussian $\mathbb{P}$, and $(\diamondsuit)$ results by computing the cross-entropy between $\nu$ and $\mathbb{Q}$. After some algebraic manipulations, we obtain $\operatorname{Tr}(\Sigma_{1}X_{\epsilon})=\operatorname{Tr}(D_{\epsilon})-\frac{\epsilon d}{4}$. Moreover, we also have that
$\log\left(\left(\tfrac{\epsilon}{2}\right)^{d}|X_{\epsilon}|\right)=\log\left(\left(\tfrac{2}{\epsilon}\right)^{d}|X_{\epsilon}^{-1}\Sigma_{2}|\right)-\log|\Sigma_{2}|,$
and, using the relationships in [7, Proposition 2.1], that
$\log|X_{\epsilon}^{-1}\Sigma_{2}|$ $=\log|X_{\epsilon}\Sigma_{1}+\tfrac{\epsilon}{2}I|$
$=\log|\Sigma_{1}^{-\frac{1}{2}}D_{\epsilon}\Sigma_{1}^{\frac{1}{2}}+\tfrac{\epsilon}{4}I|=\log|D_{\epsilon}+\tfrac{\epsilon}{4}I|.$
Therefore, we conclude that
$(\spadesuit)+(\clubsuit)+(\diamondsuit)=\ \operatorname{Tr}(\Sigma_{1})+\operatorname{Tr}(\Sigma_{2})-2\operatorname{Tr}(D_{\epsilon})+$
$\frac{\epsilon}{2}\operatorname{Tr}(\Sigma^{-1}\Sigma_{2})+\frac{\epsilon}{2}\log\frac{|\Sigma|}{|\Sigma_{2}|}+\frac{\epsilon}{2}\log\!\left(\left(\frac{2}{\epsilon}\right)^{d}\left|D_{\epsilon}+\frac{\epsilon}{4}I\right|\right).$
The claim follows from the nonnegativity of $\mathrm{KL}(\gamma\|\gamma_{0})$ and the arbitrariness of $\gamma$. ∎
Proposition 3 shows that the Sinkhorn discrepancy can be lower bounded by discarding all distributional information except for the covariances. Hence, a valid outer approximation for the set $\mathcal{S}$ is given by the entropy-regularized Gelbrich set (13), as formalized in the following Corollary.
Corollary 1:
For any regularization parameter $\epsilon\geq 0$ and radius $\rho\geq 0$, it holds that the Sinkhorn ambiguity set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho\}$ is always contained in the entropy-regularized Gelbrich set $\{\mathbb{Q}\in\mathcal{P}_{0}(\mathbb{R}^{d}):\mathbb{E}_{\mathbb{Q}}[zz^{\top}]=\Sigma_{2}\in\mathbb{S}^{d}_{++},\ G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho\}$.
By Proposition 3, we have that $W_{\epsilon}(\mathbb{P},\mathbb{Q})\geq G_{\epsilon}(\Sigma_{1},\Sigma_{2})$. Hence, if $W_{\epsilon}(\mathbb{P},\mathbb{Q})\leq\rho$, then $G_{\epsilon}(\Sigma_{1},\Sigma_{2})\leq\rho$, which completes the proof. ∎ By restricting to linear policies, we finally construct an upper bound for (7) as
$\overline{p}^{\star}=\left\{\begin{aligned} &\min_{\bf U,q}\ \max_{\mathbb{P}\in\mathcal{G}}\ \mathbb{E}_{\mathbb{P}}\left[\bf u^{\top}Ru+x^{\top}Qx\right]\\ &\text{ s.t. }\ \mathbf{U}\in\mathcal{U}^{\text{lin}},\ \bf{u=U(Dw+v)+q},\\ &\quad\quad\,\,\bf{x=Hu+Gw}.\end{aligned}\right$. (14)
As we enlarged nature’s subproblem feasible set, and at the same time shrank the possible control laws that the designer can select, we have that $\overline{p}^{\star}\geq p^{\star}$. In the next proposition, we rewrite (14) as a finite-dimensional optimization problem.
Proposition 4:
The upper bound (14) to the primal problem (5) is equivalent to the finite-dimensional program
$\overline{p}^{\star}=\min_{\begin{subarray}{c}\mathbf{U}\in\mathcal{U}^{\text{lin}}\\ \mathbf{q}\in\mathbb{R}^{mT}\end{subarray}}\ \max_{\begin{subarray}{c}\mathbf{W}\in\mathcal{G}_{W}\\ \mathbf{V}\in\mathcal{G}_{V}\end{subarray}}\ \operatorname{Tr}\left(\mathbf{F}_{1}\mathbf{W}+\mathbf{F}_{2}\mathbf{V}\right)+\mathbf{q}^{\top}\mathbf{M}\mathbf{q}$ (15)
where $\mathbf{F}_{1},\mathbf{F}_{2},\mathcal{G}_{W},\mathcal{G}_{V}$ are defined as in Proposition 2.
The proof follows the same steps as the proof of [28, Proposition 3.4]. In the same way as in Proposition 2, we first substitute $\bf{x=Hu+Gw}$ and $\bf{u=U(Dw+v)+q}$ in the objective of (14). Then, for any $\mathbb{P}\in\mathcal{G}$, we rewrite the previous expectation as a trace in terms of the covariance matrices $\mathbf{W}=\mathbb{E}_{\mathbb{P}}[\bf ww^{\top}]$ and $\mathbf{V}=\mathbb{E}_{\mathbb{P}}[\bf vv^{\top}]$, and conclude by replacing $\mathcal{G}$ with (11) and (12). ∎
We conclude this section by observing that (10) and (15) are dual to each other, in the sense that one can be obtained from the other by swapping the order of optimization.
IV-D Existence of a Nash equilibrium
In this section, we show that strong duality holds between (10) and (15), and hence that the Sinkhorn DR LQG admits a Nash equilibrium in the class of linear policies. Before proving these results, we present a technical lemma that characterizes structural and topological properties of $\mathcal{G}_{W}$ and $\mathcal{G}_{V}$.
Lemma 1:
Given $\hat{\Sigma}\in\mathbb{S}^{d}_{++}$, $\rho\geq 0$ and $\epsilon\geq 0$ and finite, the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{++}:G_{\epsilon}(\hat{\Sigma},M)\leq\rho\}$ is convex and compact.
Convexity: the map $M\rightarrow G_{\epsilon}(\hat{\Sigma},M)$ is convex because sum of convex functions. Indeed,
$G_{\epsilon}(\hat{\Sigma},M)\!=\!\underbrace{\operatorname{Tr}(\hat{\Sigma}\!+\!M\!-\!2D_{\epsilon})\!+\!\tfrac{\epsilon}{2}\left(\log\!\left(\!\left(\tfrac{2}{\epsilon}\right)^{d}\!\left|D_{\epsilon}\!+\!\tfrac{\epsilon}{4}I\right|\!\right)\right.}_{(\heartsuit)}\!\left.\!+\operatorname{Tr}\!\left(\Sigma^{-1}M\right)\!+\!\log\tfrac{|\Sigma|}{|M|}\right)\!,$
and $(\heartsuit)$ is convex because of [18, Proposition 6], while the trace is linear (hence convex) and the negative log-determinant is convex. This implies that $\mathcal{D}$ is convex because it is the level set of a convex function [26, Proposition 2.7].
Compactness: we want to show that $\mathcal{D}$ is closed and bounded. To this end, let the function $f:\mathbb{S}^{d}_{++}\rightarrow\mathbb{R}$ be defined as
$f(M)=\operatorname{Tr}(M-2D_{\epsilon})+\frac{\epsilon}{2}\left(\log|D_{\epsilon}+\frac{\epsilon}{4}I|+\operatorname{Tr}(\Sigma^{-1}M)-\log|M|\right)$.
This function involves affine transformations of $M$ along with continuous transformations on $\mathbb{S}^{d}_{++}$ such as matrix square-root, trace and log-determinant. Hence, $f(\cdot)$ is continuous and the set $\mathcal{D}=\{M\in\mathbb{S}^{d}_{++}:f(M)\leq\tilde{\rho}\}$ with $\tilde{\rho}=\rho-\operatorname{Tr}(\hat{\Sigma})+\frac{\epsilon}{2}(\log(\frac{\epsilon}{2})^{d}-\log|\Sigma|)$ is closed because it is the lower level set of a continuous function [26, Theorem 1.6].
To show boundedness, we proceed by contradiction and assume that $\sup_{M\in\mathcal{D}}\ \lambda_{\max}(M)=+\infty$. We construct a lower bound for $f(M)$ by bounding each addend separately. The term $\operatorname{Tr}(M)$ can be lower bounded by $\lambda_{\max}(M)$. Since $\|\hat{\Sigma}^{\frac{1}{2}}M\hat{\Sigma}^{\frac{1}{2}}\|_{2}\leq\|\hat{\Sigma}^{\frac{1}{2}}\|_{2}^{2}\|M\|_{2}=\lambda_{\max}(\hat{\Sigma})\lambda_{\max}(M)$ by the submultiplicativity property of the operator norm, $\operatorname{Tr}(D_{\epsilon})=\sum_{i=1}^{d}\sqrt{\lambda_{i}(\hat{\Sigma}^{\frac{1}{2}}M\hat{\Sigma}^{\frac{1}{2}})+\frac{\epsilon^{2}}{16}}\leq d\sqrt{\lambda_{\max}(\hat{\Sigma})\lambda_{\max}(M)+\frac{\epsilon^{2}}{16}}$, and we obtain a lower bound for the second addend. By definition of $D_{\epsilon}$ in Definition 3, since $\hat{\Sigma}^{\frac{1}{2}}M\hat{\Sigma}^{\frac{1}{2}}\succ 0$, we have that $\lambda_{i}(D_{\epsilon})>\frac{\epsilon}{4}\,\forall i$. Therefore, $\left|D_{\epsilon}+\frac{\epsilon}{4}I\right|\geq\left(\frac{\epsilon}{2}\right)^{d}$ and we can lower bound the third addend by $\frac{\epsilon d}{2}\log\left(\frac{\epsilon}{2}\right)$. The fourth addend can be bounded as $\operatorname{Tr}(\Sigma^{-1}M)\geq\frac{\operatorname{Tr}(M)}{\lambda_{\max}(\Sigma)}\geq\frac{\lambda_{\max}(M)}{\lambda_{\max}(\Sigma)}$. Finally, $\log|M|=\sum_{i=1}^{d}\log(\lambda_{i}(M))\leq d\log(\lambda_{\max}(M))$ and $\log\frac{|\Sigma|}{|M|}$ is lower-bounded by $d\log|\Sigma|\log\lambda_{\max}(M)$. Putting everything together we get
$f(M)\geq\ \lambda_{\max}(M)-2d\sqrt{\lambda_{\max}(\hat{\Sigma})\lambda_{\max}(M)+\frac{\epsilon^{2}}{16}}$
$+\frac{\epsilon d}{2}\log\left(\frac{\epsilon}{2}\right)+\frac{\epsilon}{2}\frac{\lambda_{\max}(M)}{\lambda_{\max}(\hat{\Sigma})}-\frac{\epsilon d}{2}\log|\Sigma|\log\left(\lambda_{\max}(M)\right).$
As $\lambda_{\max}(M)\rightarrow+\infty$, the linear terms in $\lambda_{\max}(M)$ dominate the other ones. Consequently, the RHS is unbounded when choosing $M$ such that $\lambda_{\max}(M)=+\infty$. Therefore, since $\tilde{\rho}$ is finite when so is $\epsilon$, we contradict the fact that any $M\in\mathcal{D}$ satisfies $f(M)\leq\tilde{\rho}$. Hence, $\mathcal{D}$ is bounded. ∎ Lemma 1 is key, as it enables the use of Sion’s minimax theorem [27, Theorem 3.4] to prove the existence of a Nash equilibrium for (5). This is formally stated in the next theorem.
Theorem 1:
The following results hold:
  1. 1.
    The optimal values $\overline{p}^{\star}$ of (9) and $\underline{d}^{\star}$ of (14) coincide;
  2. 2.
    The optimal values $p^{\star}$ of (7) and $d^{\star}$ of (8) coincide;
  3. 3.
    There exist $\mathbf{U}^{\star}\in\mathcal{U}^{\text{lin}}$ and $\mathbf{q}^{\star}\in\mathbb{R}^{mT}$ such that the DR LQG problem (7) is solved by $\bf u=q^{\star}+U^{\star}y$;
  4. 4.
    The dual DR problem (8) is solved by a Gaussian distribution $\mathbb{P}^{\star}\in\mathcal{S_{N}}$.
By Lemma 1 the sets $\mathcal{G}_{W}$ and $\mathcal{G}_{V}$ are compact and convex. The trace is linear, hence the objective function is concave in $\mathbf{W}$ and $\mathbf{V}$. Moreover, since $\mathbf{Q},\mathbf{R},\mathbf{M}\succeq 0$, it is convex in $\mathbf{U}$ and $\mathbf{q}$. Therefore, we can apply Sion’s minimax theorem [27, Theorem 3.4] to show that strong duality holds, proving the first point of the theorem. By strong duality, the chain of inequalities $\underline{d}^{\star}\leq d^{\star}\leq p^{\star}\leq\overline{p}^{\star}$ collapses to equalities proving the second point. The equality $p^{\star}=\overline{p}^{\star}$ implies that (7) is solved by a linear causal policy of the purified observations. However, as pointed out in Section IV-A, any causal controller that is linear in the purified outputs $\bm{\eta}$ can be also expressed as a causal linear feedback in the measurements $\bf y$. This proves the third point. We finish the proof noticing that the last point of the theorem follows from the identity $\underline{d}^{\star}=d^{\star}$. ∎ Theorem 1 implies that (5) is solved by an LQG with worst-case covariance matrices given by (10). We report a reformulation of (10) as finite-dimensional conic program in Proposition 5 in the Appendix. We conclude this section by observing that, when $\epsilon\rightarrow\infty$, the set $\mathcal{S}$ either becomes empty or reduces to the singleton $\nu$ depending on whether $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)$ exceeds $\rho$ or not, see [6, Proposition 1]. In particular, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)\leq\rho$, we still retain global optimality of linear policies as $\mathcal{S}=\{\nu\}$, and $\nu$ is Gaussian. Instead, when $\operatorname{Tr}(\hat{\Sigma})+\operatorname{Tr}(\Sigma)>\rho$, the Sinkhorn DR LQG problem (5) becomes infeasible.
V Numerical Experiments
In this section, we present numerical simulations to showcase the advantages of robustifying against distributional uncertainty. [^4] All our experiments were run on an M3 Pro CPU machine with 36GB RAM. All SDP problems were modeled in Matlab 2023a using Yalmip and solved with MOSEK. Our source code is publicly available at https://github.com/DecodEPFL/Optimality_Sinkhorn.git. For our experiments, we consider the open-loop unstable discrete-time linear dynamical system given by
$x_{t+1}=\begin{bmatrix}1.1&0.1\\ 0&1.1\end{bmatrix}x_{t}+\begin{bmatrix}1\\ 1\end{bmatrix}u_{t}+w_{t},\quad y_{t}=x_{t}+v_{t}\,$,
with cost matrices $Q_{t}=I_{2}$ and $R_{t}=1$ at all times, control horizon $T=25$, and nominal covariances $\hat{X}_{0}=I_{2},\hat{W}_{t}=I_{2},\hat{V}_{t}=0.01I_{2},\ \forall t\in[T]$. The reference covariance was picked $\Sigma=I_{2}$.
In Fig. 1, we benchmark the classical LQG controller designed based on the nominal covariances against the Sinkhorn DR LQG policy obtained by solving (5) using the conic program in Proposition 5 with radii $\rho_{x_{0}}=20,\rho_{w_{t}}=0.2,\rho_{v_{t}}=20$ and regularization parameter $\epsilon=0.01$. Specifically, we compare the performance of these controllers on $5000$ realizations of the exogenous disturbances drawn from the nominal Gaussian distribution (on the left) and from the respective nature’s adversarial choice of distribution $\mathbb{P}$ within $\mathcal{S}$ (on the right). As expected, we observe that the Sinkhorn DR LQG policy incurs a slightly higher average cost when the true distribution corresponds to the nominal one. Conversely, when the noise distributions are selected adversarially within $\mathcal{S}$, we observe that the proposed Sinkhorn DR LQG policy achieves a lower average cost. These results validate our design and highlight fundamental tradeoff between nominal performance and distributional robustness.
Figure 1: Comparison between the control cost incurred by the nominal LQG controller (red histograms, in the background) and the proposed Sinkhorn DR LQG policy (green histograms, in the foreground) over 50005000 disturbance realizations drawn from the nominal distribution (on the left) and the respective worst-case distribution in the Sinkhorn ambiguity set (on the right). Dotted vertical lines represent theoretical mean values.
VI Conclusion
In this work, we studied a DR generalization of classical LQG control where the noise distributions are unknown and belong to entropy-regularized Wasserstein or Sinkhorn ambiguity sets. We proved that, despite robustifying the objective in the probability space, nature’s distributions retain a Gaussian form and hence linear policies remain globally optimal. We validated the effectiveness of our Sinkhorn DR LQG policy through numerical simulations showing improved robustness compared to classical LQG design.
Appendix A Appendix
We report below expressions for the stacked cost matrices $\mathbf{Q}\in\mathbb{S}_{+}^{d(T+1)}$ and $\mathbf{R}\in\mathbb{S}_{++}^{mT}$, and for the matrices $\mathbf{C}\in\mathbb{R}^{pT\times d(T+1)}$, $\mathbf{G}\in\mathbb{R}^{d(T+1)\times d(T+1)}$, and $\mathbf{H}\in\mathbb{R}^{d(T+1)\times mT}$ encoding the system linear dynamics and measurement equations. Specifically, we have $\mathbf{Q}=\operatorname{blkdiag}(Q_{0},\dots,Q_{T})$, $\mathbf{R}=\operatorname{blkdiag}(R_{0},\dots,R_{T-1})$ and
$\small\mathbf{C}=\begin{bmatrix}C_{0}&0&&&\\ &C_{1}&\ddots&&\\ &&\ddots&\ddots&\\ &&&C_{T-1}&0\end{bmatrix}\,,\quad\mathbf{G}=\begin{bmatrix}A^{0}_{0}&&&\\ A_{0}^{1}&A_{1}^{1}&&\\ \vdots&&\ddots&\\ A^{T}_{0}&A^{T}_{1}&\dots&A^{T}_{T}\end{bmatrix}\,$,}
$\small\mathbf{H}=\begin{bmatrix}0&&&&\\ A^{1}_{1}B_{0}&0&&&\\ A_{1}^{2}B_{0}&A_{2}^{2}B_{1}&0&&\\ \vdots&&&\ddots&\\ \vdots&&&&0\\ A^{T}_{1}B_{0}&A^{T}_{2}B_{1}&\dots&\dots&A^{T}_{T}B_{T-1}\end{bmatrix}$,}
where $A^{t}_{s}=\prod\limits_{k=s}^{t-1}A_{k}$ for any $s<t$ and $A_{s}^{t}=I$ for $s=t$.
Proposition 5:
The optimization problem (10) is equivalent to the conic program
$\max\$$\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)},\ \mathbf{V}\in\mathbb{S}_{++}^{pT},$
$E_{x_{0}}\in\mathbb{S}_{+}^{d},\ E_{w_{t}}\in\mathbb{S}_{+}^{d},\ E_{v_{t}}\in\mathbb{S}_{+}^{p},\ \forall t\in[T],$
$\begin{aligned} &\operatorname{Tr}\!\left(\!W_{0}\!-\!2E_{x_{0}}\!+\!\frac{\epsilon}{2}\Sigma^{-1}W_{0}\!\right)\!-\!\frac{\epsilon}{2}\log\left|E_{x_{0}}\!-\frac{\epsilon}{4}I\right|\leq\rho_{x_{0}}\\ &-\operatorname{Tr}(\hat{X}_{0})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{X}_{0}\Sigma|\right),\end{aligned}$
$\begin{aligned} &\operatorname{Tr}\left(\!W_{t+1}\!\!-\!2E_{w_{t}}\!\!+\!\frac{\epsilon}{2}\Sigma^{-1}W_{t+1}\!\right)\!\!-\!\frac{\epsilon}{2}\!\log\!\left|E_{w_{t}}\!\!-\!\frac{\epsilon}{4}I\right|\!\!\leq\!\rho_{w_{t}}\\ &-\operatorname{Tr}(\hat{W}_{t})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{W}_{t}\Sigma|\right),\ \forall t\in[T]\end{aligned}$
$\begin{aligned} &\operatorname{Tr}\left(V_{t}\!-\!2E_{v_{t}}\!+\!\frac{\epsilon}{2}\Sigma^{-1}V_{t}\right)\!-\!\frac{\epsilon}{2}\log\left|E_{v_{t}}\!-\!\frac{\epsilon}{4}I\right|\leq\rho_{v_{t}}\\ &-\operatorname{Tr}(\hat{V}_{t})+\frac{\epsilon}{2}\left(d\log\left(\frac{\epsilon}{2}\right)-\log|\hat{V}_{t}\Sigma|\right),\ \forall t\in[T]\end{aligned}$
$\begin{bmatrix}\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{x_{0}}\\ E_{x_{0}}&I_{n}\end{bmatrix}\succeq 0,$
$\begin{bmatrix}\hat{W}_{t}^{1/2}W_{t+1}\hat{W}_{t}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{w_{t}}\\ E_{w_{t}}&I_{n}\end{bmatrix}\succeq 0,\ \forall t\in[T],$
$\begin{bmatrix}\hat{V}_{t}^{1/2}V_{t}\hat{V}_{t}^{1/2}+\frac{\epsilon^{2}}{16}I_{p}&E_{v_{t}}\\ E_{v_{t}}&I_{p}\end{bmatrix}\succeq 0,\ \forall t\in[T],$
$\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!0\,.$
Here, $\mathcal{M}$ denotes the set of all strictly upper block triangular matrices of the form
$\Gamma=\begin{bmatrix}0&\Gamma_{1,2}&\Gamma_{1,3}&\cdots&\Gamma_{1,T}\\ &0&\Gamma_{2,3}&\cdots&\Gamma_{2,T}\\ &&\ddots&\ddots&\vdots\\ &&&0&\Gamma_{T-1,T}\\ &&&&0\end{bmatrix}\in\mathbb{R}^{Tm\times Tp}\,$,
where $\Gamma_{t,s}\in\mathbb{R}^{m\times p}$ for every $t,s\in\mathbb{Z}$ with $1\leq t<s\leq T$.
The proof relies on dualizing the inner minimization problem in (10). Strong duality holds because the primal problem is feasible and involves only equality constraints, hence any feasible point is in fact a Slater point.
In the following we use $\Gamma\in\mathcal{M}$ to denote the Lagrange multiplier of the constraint $\bf U\in\mathcal{U}^{\text{lin}}$, which requires all blocks of $\bf U$ above the main diagonal to vanish. The Lagrangian function of the inner minimization problem in (10) is
$\mathcal{L}(\mathbf{q},\mathbf{U},\Gamma)=\operatorname{Tr}\left(\mathbf{F}_{1}\mathbf{W}+\mathbf{F}_{2}\mathbf{V}\right)+\mathbf{q}^{\top}\mathbf{M}\mathbf{q}+\operatorname{Tr}(\mathbf{U}\Gamma^{\top})$
Recall now that $\mathbf{R}\succ 0$ and $\mathbf{Q}\succeq 0$, and thus $\mathbf{M}\succ 0$. Consequently, $\mathcal{L}$ is minimized by $\mathbf{q}^{\star}=0$ for any fixed $\bf U$ and $\Gamma$. In addition, the partial gradient of $\mathcal{L}$ with respect to $\bf U$ is given by
$\frac{\partial\mathcal{L}}{\partial\bf U}=2\mathbf{MUDWD}^{\top}+2\mathbf{MUV}+2\mathbf{H^{\top}QGWD^{\top}}+\Gamma\,.$
Recall also that $\mathbf{V}\in\mathcal{G}_{V}$ is strictly positive, which implies that $\mathbf{DWD^{\top}}+\mathbf{V}\succ 0$ is invertible. As we already know that $\mathbf{M}\succ 0$ is invertible as well, $\mathcal{L}$ is minimized by
$\mathbf{U}^{\star}=-\mathbf{M}^{-1}(\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)(\mathbf{DWD^{\top}+V})^{-1}$
for any fixed $\Gamma$.
Substituting both $\bf q^{\star}$ and $\bf U^{\star}$ into $\mathcal{L}$ yields the dual objective function
$g(\Gamma)=\mathcal{L}(\mathbf{q^{\star},U}^{\star},\Gamma)=\operatorname{Tr}(\mathbf{G^{\top}QGW})+$
[…]
The dual of the inner minimization problem in (10) is thus given by $\max_{\Gamma\in\mathcal{M}}g(\Gamma)$. To linearize the dual function $g(\Gamma)$, we introduce the auxiliary variable $F\in\mathbb{S}_{+}^{mT}$ subject to the matrix inequality
[…]
We can reformulate the previous inequality as an LMI using the Schur complement. Then, the dual problem rewrites as
$\max$ $\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$ (16)
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},$
$\!\!\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!\!0\,.$
Substituting the strong dual (16) in the inner minimization in (10), we obtain
$\max\ \$$\operatorname{Tr}(\mathbf{G^{\top}QGW})-\operatorname{Tr}\big(\mathbf{M}^{-1}F\big)$
s.t. $\Gamma\in\mathcal{M},\ F\in\mathbb{S}_{+}^{mT},\mathbf{W}\in\mathbb{S}_{++}^{d(T+1)},\ \mathbf{V}\in\mathbb{S}_{++}^{pT},$
$\!\!\begin{bmatrix}F&\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2\\[1.0pt] (\mathbf{H^{\top}QGWD^{\top}}+\Gamma/2)^{\top}&\mathbf{DWD^{\top}+V}\end{bmatrix}\!\!\succeq\!0,$
$G_{\epsilon}(\hat{X}_{0},W_{0})\leq\rho_{x_{0}},$
$G_{\epsilon}(\hat{W}_{t+1},W_{t+1})\leq\rho_{w_{t}},\ G_{\epsilon}(\hat{V}_{t},V_{t})\leq\rho_{v_{t}}\ \forall t\in[T]\,.$
We conclude the proof by reformulating the entropy-regularized Gelbrich constraints. For simplicity, consider the constraint $G_{\epsilon}(\hat{X}_{0},W_{0})\leq\rho_{x_{0}}$; the reasoning applies mutatis mutandis to the others. Consider an auxiliary variable $E_{x_{0}}\in\mathbb{S}^{n}_{+}$ subject to the matrix inequality $E_{x_{0}}^{2}\preceq\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}$. This inequality can be recast as $E_{x_{0}}\preceq(\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n})^{\frac{1}{2}}$. With this auxiliary variable, we can equivalently express the nonlinear inequality constraint as
[…]
If the constraint is satisfied with $E_{x_{0}}$ it will be also with the original variable since the trace and the log-determinant are monotone operators. Finally, with a Schur complement argument the inequality $E_{x_{0}}^{2}\preceq\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}$ can be equivalently rewritten as
$\begin{bmatrix}\hat{X}_{0}^{1/2}W_{0}\hat{X}_{0}^{1/2}+\frac{\epsilon^{2}}{16}I_{n}&E_{x_{0}}\\ E_{x_{0}}&I_{n}\end{bmatrix}\succeq 0\,.$